Fading display 475


I have a strange issue with my 475 which otherwise works normally. On switching on, the display appears, but if I adjust the beam intensity knob either way the trace slowly fades. Not to reappear unless I turn off and then back on. This can repeat a few times until the trace remains stable and I can adjust the intensity normally. The intensity potentiometer tests normal. I thought it might be a poor connection at the base of the CRT, but I have cleaned the contacts and applied switch cleaner.
I sense there is a poor contact somewhere, but don’t know where to look.


The behavior sounds like something that I experienced within the last few months, but I don't have a good characterization for the symptoms. The good news is that I think what I experienced was "operator error" rather than some fault in the instrument. What I was seeing was that I would get a stable triggered display, then it would lose trigger and the trace would slowly fade away. Once the trigger was lost neither beam intensity controls nor beam finder would bring up anything on the CRT, because there was nothing to show, of course.

I think that the behavior I experienced was with the scope in NORM trigger mode (because it must have been, since any other mode would have still generated a periodic sweep). I don't understand how I was getting this slow fade out of the trace, however. I also don't recall which instrument I was experiencing this problem with, though I doubt that it was with a 475 (it was probably a 2215A, 2252, 2465, or possibly a 5103 or 7603).

-- Jeff Dutky

Mark Vincent


Clean the beam finder switch. That is in the intensity circuit. These switches will cause problems when dirty because they are used infrequently. This could be the cause of your problem. Try this first.

If the high voltage circuit area is dirty, clean it. The dust can/will cause leakage paths that cause problems

Q1338 is a 151-0367-00 (a 151-0402-00 is the same transistor with a different number on it and has the same leakage problem) type that I have found to be leaky from C to E as a diode. All I find bad I replace with KSP10BU. These have a BEC pinout. This model has some of these in it. The only ones I find that are not leaky are the ones with a light blue tint on the face of the transistor.

The 1mfd filter, C1374, from the bias control I replace with a 200/250V film type. The ESR on the original electrolytics I have found to be high to open. This value is not critical. A 1,5mfd, for example, will work fine.

R1378, 22meg 1/4W carbon, I replace with RT37 types that are 1/2W 1%. Testing these I found mine to be well high in resistance, 30-50meg. When high, the trace is bright then drops back down as it warms for 10-20 minutes.

Check the 9V zener, VR1416, in the power supply. It is likely open. I have found when this is open, the voltages drift with temperature. Once I replaced the zener with a BZX85B9V1, the voltages were stable. In this model or A version, I changed the 5620 ohm 1/2W resistor to a 5600 ohm 2W 1% resistor. The original resistor was dark in the middle of the body and out of tolerance.

What I said I did in mine, you can do in yours if you want. I stated what I found wrong in mine and corrected.



Thanks, that gives me some leads to try.


Thanks, I'll check the triggering.